OpenStudy (anonymous):

What is the equation of the line passing through (1, 7) and parallel to the line y = –x + 5 in standard form?

OpenStudy (anonymous):

y=-x+6

OpenStudy (anonymous):

medal please

OpenStudy (anonymous):

you're wrong @ivan514

OpenStudy (nali):

since it has to be parallel use the same slope and then just find b by plugging in (1,7) in to y=mx+b

OpenStudy (anonymous):

x-y=8?

OpenStudy (nali):

y=mx+b 7=(-1) (1) + b 7=-1+b b=8

OpenStudy (anonymous):

what is that in standard form

OpenStudy (nali):

y=-x+8 you mean ax+by+c=0 then it is this x+y-8=0 when you move the x and the 8 to the other side